# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v5

@@ G.1.1 Primary ligand RMSD results
In Figures 20, 21, and 22, we report the (binding site-superimposed) ligand RMSD values of each baseline method across the primary ligand Astex Diverse, PoseBusters Benchmark, and DockGen datasets, with molecular dynamics (MD)-based structural relaxation applied post-hoc. Overall, these figures demonstrate that AF3 and Chai-1 achieve the tightest RMSD distributions, except for single-sequence AF3 which occasionally produces catastrophic prediction errors by targeting incorrect PLI binding pockets. Further, these results show that MD-based relaxation generally does not modify the RMSD distribution of most baseline methods, except for DynamicBind and RFAA for which neither seem to benefit from such post-hoc optimizations.
@@ G.2.3 All single-ligand results
Next, Figures 26, 27, 28, and 29 display each method’s single-ligand CASP15 docking success rates, PB-Valid rates, docking RMSD, and docking lDDT-PLI distributions, respectively. In summary, we can make a few respective observations from these plots. (1) AF3 achieves the highest structural accuracy for this subset of targets yet is challenged in its PLIF-WM rate by conventional and DL co-folding baseline methods such as AutoDock Vina and NeuralPLexer. (2) Even though most are positionally incorrect, structurally and chemically speaking, the majority of AutoDock Vina and DiffDock-L’s predictions are valid according to the PoseBusters software suite, whereas fewer of AF3’s predictions are. (3) AutoDock Vina, DiffDock-L, NeuralPLexer, and AF3 yield notably lower RMSD distributions than all other baseline methods (including all single-sequence DL co-folding variants). (4) Only AF3 and AutoDock Vina produce a reasonable range of lDDT-PLI scores for these single-ligand targets.
@@ G.2.4 Single and multi-ligand results for public targets
Lastly, for completeness and reproducibility, Figures 30, 31, 32, and 33 present corresponding multi-ligand results for the public CASP15 targets, whereas Figures 34, 35, 36, and 37 report corresponding single-ligand results for the public CASP15 targets. Overall, we observe marginal differences between the full and public CASP15 target results for multi-ligand complexes, since once again AF3 achieves top results overall in the context of multi-ligands. However, we notice more striking performance drops between the full and public single-ligand CASP15 target results, suggesting that some of the private single-ligand complexes are easier prediction targets than most of the publicly available single-ligand complexes. In short, we find that AutoDock Vina consistently performs best in this single-ligand setting.

### v6

@@ G.1.1 Primary ligand RMSD results
In Figures 20, 21, and 22, we report the (binding site-superimposed) ligand RMSD values of each baseline method across the primary ligand Astex Diverse, PoseBusters Benchmark, and DockGen datasets, with molecular dynamics (MD)-based structural relaxation applied post-hoc. Overall, these figures demonstrate that Chai-1, Boltz-1, and AF3 achieve the tightest RMSD distributions, except for single-sequence AF3 which occasionally produces catastrophic prediction errors by targeting incorrect PLI binding pockets. Further, these results show that MD-based relaxation generally does not modify the RMSD distribution of most baseline methods, except for DynamicBind and RFAA for which neither seem to benefit from such post-hoc optimizations.
@@ G.2.3 All single-ligand results
Next, Figures 26, 27, 28, and 29 display each method’s single-ligand CASP15 docking success rates, PB-Valid rates, docking RMSD, and docking lDDT-PLI distributions, respectively. In summary, we can make a few respective observations from these plots. (1) Interestingly, Boltz-1 (with and without MSAs) achieves the highest structural and interaction modeling accuracy compared to AF3 and similar co-folding methods. (2) Even though most are positionally incorrect, structurally and chemically speaking, the majority of Boltz-1, AutoDock Vina, and DiffDock-L’s predictions are valid according to the PoseBusters software suite, whereas few of AF3’s predictions are. (3) Boltz-1 and AF3 yield notably lower RMSD distributions than all other baseline methods (including the similar DL co-folding method Chai-1). (4) Only AutoDock Vina, Boltz-1, and AF3 produce a reasonable range of lDDT-PLI scores for these single-ligand targets.
@@ G.2.4 Single and multi-ligand results for public targets
Lastly, for completeness and reproducibility, Figures 30, 31, 32, and 33 present corresponding multi-ligand results for the public CASP15 targets, whereas Figures 34, 35, 36, and 37 report corresponding single-ligand results for the public CASP15 targets. Overall, we observe marginal differences between the full and public CASP15 target results for multi-ligand complexes, since once again AF3 achieves top results overall in the context of multi-ligands. However, for most methods, we notice more striking performance drops between the full and public single-ligand CASP15 target results, suggesting that some of the private single-ligand complexes are easier prediction targets than most of the publicly available single-ligand complexes. In short, we find that Boltz-1 consistently performs best in this single-ligand setting.